radomes.openskiesproject.org
avataytay03
avataytay03 avataytay03
  • 27-04-2015
  • Mathematics
Answered



7) Jan has two spinners. One has 3 equal sections labeled A, B, and C. The other has 4 equal sections labeled 1, 2, 3, and 4. She spins both spinners. What is the probability that she gets A3?

Verified Answer :

JaySL JaySL
  • 27-04-2015

One out of three results on the first spinner is A, and one out of four on the second is 3.  Thus, the overall probability is 1/3*1/4=1/12.

Answer Link

Related Questions

The muscles in the stomach aid in digestion by
what is the name of the large island body of water bordered by Iran azerbaijan russia kazakhstan and Turkmenistan
use the term technolgyscientific law and scientific theory in a complete sentence
Which describes what happens when a balloon is rubbed with a wool cloth
The formula of the area of a triangle is A12bh Solve the formula for h
During the colonial era in America slave codes
describe what happens when an egg is fertilized in a female
simplified form of 354756
How did the beliefs of Booker T Washington and WEB Du Bois differ A Washington fought for economic equality Du Bois fought for social equality B Washington foug
Our healthcare may improve if everyone ate healthier Which of these will make the sentence conditional in mood ate may healthier improve